Overhead Budget Johnston Company cleans and applies powder coat paint to metal items on a job-order basis. Jonnston has budgeted the following amounts for various overhead categories in the coming year Supplies Gas Indirect labor Supervision Depreciation on equipment Depreciation on the buliding Rental of special equipment Electricity (for lighting, heating, and air conditioning) Telephone Landscaping service Other overhead $218,000 54,000 172,000 75,000 54,000 47,000 14,000 27,500 4,900 1,200 54,000 In the coming year, Johnston expects to powder coat 160,000 units. Each unit takes 1:2 direct labor hours. Johnston has found that supplies and gas (used to run the drying ovens-all units pass through the drying ovens after powder coat paint is applied) tend to vary with the number of units produced. All other overhead categories are considered to be fixed. Required: 1, Calculate the number of direct labor hours Johnston must budget for the coming year. Calculate the vanable overhead rate. Calculate the total fixed overhead for the coming year. When required, round your answers to the nearest cent and use the rounded answer in subsecuent recuirements 1.2 X Direct labor hours Variable overhead rate Total fixed overhead per direct labor hou

media%2Fe53%2Fe532e021-a01e-4fc5-911b-e2

media%2F13d%2F13da8727-9625-4f7e-bbbc-9f

media%2F8bb%2F8bb3bfd3-14d0-4d97-abd7-1e
Answer #1

1)Direct labor hours = 1.2*160000= 192000 DLH

2)Total variable overhead = supplies + gas

            = 218000+54000

            = 272000

variable overhead rate= Total variable overhead rate /Direct labor hours

       = 272000/192000

      = $ 1.42 per DLH

Fixed overhead = 172000+75000+54000+47000+14000+27500+4900+1200+54000

               = 449600

2)

Budgeted direct labor hours 192000
variable overhead rate 1.42
Budgeted variable overhead [192000*1.42] 272640
Budgeted fixed overhead 449600
Total budgeted overhead 722240

Fixed overhead rate = 449600/192000= $ 2.34 per DLH

Total overhead rate= 722240/192000= $ 3.76 per DLH

3)New budgeted Direct labor hours = new units *hours per unit

                  = 158000*1.2

                   = 189600 DLH